# LoRA vs Full Fine-tuning: An Illusion of Equivalence

### v1

@@ 5 Why Do Intruder Dimensions Exist?
Differences in the update rule: As described in Appendix D, LoRA and full fine-tuning have characteristically different update rules, even for the same training examples. We highlight that LoRA uses a larger learning rate and has gradients projected into a low-rank space (Hao et al., 2024), leading to conditions similar to the toy example above.
@@
Figure 10: Impact of only tuning B on the number of intruder dimensions. We randomly initialize A such that it has singular values of 1, freeze it, and only train B. When we do this, we see a sharp reduction in high ranking intruder dimensions in comparison to those in normal LoRA (reported in Fig. 5(a)). Graphs for a specific dataset have the same range as Fig. 5(a) for easy comparison.
@@
Product parameterization of LoRA: Multiplying matrices together amplifies their spectral differences (their singular values) and in most cases leads to a lower effective rank. To test the impact of the product $BA$ on the introduction of intruder dimensions, we randomly initialize $A$ such that all its singular values are 1 and freeze it. We only tune $B$ and keep the rest of our fine-tuning recipe the same. Comparing this with vanilla LoRA is fair because Zhu et al. (2024) found that tuning $B$ is more impactful and important for generalization in comparison to $A$ and Hao et al. (2024) showed that only tuning $B$ effectively approximates LoRA. As we can see in Fig. 10, we see a sharp drop in the number of high ranking intruder dimensions when only tuning $B$ in comparison to the vanilla LoRA case where we train $A$ and $B$ separately, as reported in Fig. 5. This suggests that the matrix product of LoRA is an important component in the introduction of intruder dimensions because of how it amplifies the spectral differences of $B$ and $A$.
6 Conclusion
The paper describes the finding that LoRA and full fine-tuning, with equal performance on the fine-tuning task, can have solutions with very different generalization behaviors outside the fine-tuning task distribution. We found that LoRA and full fine-tuning yield models with significant differences spectral properties of their weight matrices: LoRA models often containing “intruder dimensions”, high-ranking singular vectors approximately orthogonal to the singular vectors of pre-trained weight matrices. The existence of intruder dimensions correlates with the fine-tuned model forgetting more of the pre-training distribution as well as forgetting more when trained on tasks sequentially in a continual learning setup.
Acknowledgements
We would like to thank Jacob Portes and Dan Biderman for corresponding with us, releasing their LLaMA-2 7B checkpoints for us to use, and even running another LoRA fine-tune with $r=2048$ for our work. This enabled us to study a more comprehensive range of models. We would also like to thank Leshem Chosen, Lucas Hennigen, Han Guo, and the entire Language & Intelligence lab for their helpful feedback on this work. This research was supported in part by the National Science Foundation under grant IIS-2238240.
Appendix A RoBERTa fine-tuning details
@@
We generally follow the procedure used by Hu et al. (2021). For all models, we use a linear learning rate schedule with 0.06 linear warmup ratio and train for a maximum of 5 epochs with batch size 16. We use the Adam optimizer with no weight decay and a maximum sequence length of 512. We fine-tune all linear layers besides the embedding matrix as well as all bias and LayerNorm layers to ensure fair comparison between methods. For full fine-tuning, we use a learning rate of 1e-5. For LoRA, we set $\alpha=2r$, and train for all ranks in {1, 2, 4, 8, 16, 64, 768}. We hold the “total learning rate of LoRA”, which is $\alpha*\eta$, fixed as we sweep rank such that this product always equals 2.4e-3. We train these models to equivalent accuracy on their downstream task. We fine-tune on six sequence classification tasks: sentiment analysis (Socher et al., 2013), entailment (Williams et al., 2018), duplicate identification (Wang et al., 2019), fact verification (Thorne et al., 2018), and common sense reasoning (Sap et al., 2019; Sakaguchi et al., 2021).
Appendix B Model Accuracies
@@
We provide the accuracies that our RoBERTa models achieve in Table 1 and Table 2.
@@
RoBbase Full 0.8617 0.9461 0.9146 0.6251 0.6551 0.6687
r=1 0.8615 0.9427 0.9033 0.6212 0.6305 0.6794
r=2 0.8639 0.9392 0.9053 0.6369 0.6530 0.6663
r=4 0.8615 0.9438 0.9083 0.6440 0.6633 0.6667
r=8 0.8707 0.9415 0.9079 0.6322 0.6571 0.6739
r=16 0.8666 0.9495 0.9088 0.6338 0.6679 0.6730
r=64 0.8710 0.9473 0.9073 0.6283 0.6274 0.6780
r=768 0.8690 0.9381 0.9024 0.6133 0.6274 0.6729
Table 2: Model accuracies on their given downstream task after fine-tuning for $\alpha=2r$.
Appendix C Cosine Similarity with Orthogonal Vectors that Span a Space
@@
Appendix D Derivation of Gradients
Our calculations follow a similar line to that of Hao et al. (2024).
@@ Appendix D Derivation of Gradients
Appendix E Impact of Matrix Percentage on Number of Intruder Dimensions
@@
Appendix F Plots of Impact of Dataset Size
Figure 12: (Top) Impact of cosine similarity threshold, $\epsilon$, on the number of intruder dimensions for LoRA models trained on different proportions of the MNLI dataset. (Bottom) Impact of the number of fine-tuned singular vectors we examine, $k$, on the number of intruder dimensions for LoRA models trained on different proportions of the MNLI dataset. We see that training on a larger proportion of the dataset increases the number of intruder dimensions in the model.
Appendix G Effective Rank when alpha=2r
Figure 13: Effective Rank of RoBERTa on MNLI when $\alpha=2r$. We measure for all weight types. For a specific weight type, the graph on the left shows the effective rank of all models, and the right shows the effective rank of the LoRA models only.
Appendix H LLaMA/LLaMA-2 Instruction Tuned Models
Our LLaMA-7B checkpoints were fine-tuned on the Alpaca (Taori et al., 2023) and consist of two fully fine-tuned models, one LoRA model with rank 16, and one QLoRA (Dettmers et al., 2023) model with rank 64. Our LLaMA2-7B checkpoints were fine-tuned on either Magicoder-Evol-Instruct-110K (Wei et al., 2024) or MetaMathQA (Yu et al., 2024) and consist of one fully fine-tuned model and 3-4 LoRA’ed models of different ranks for each dataset and generously provided by Biderman et al. (2024).
@@ Appendix H LLaMA/LLaMA-2 Instruction Tuned Models
LoRA-TMLR-2024/magicoder-lora-rank-2048-alpha-4096 LLaMA2-7b Magicoder
@@
LoRA-TMLR-2024/magicoder-lora-rank-16-alpha-32 LLaMA2-7b MetaMath
LoRA-TMLR-2024/magicoder-lora-rank-64-alpha-128 LLaMA2-7b MetaMath
LoRA-TMLR-2024/magicoder-lora-rank-256-alpha-512 LLaMA2-7b MetaMath
LoRA-TMLR-2024/magicoder-full-finetuning-lr-1e-05 LLaMA2-7b MetaMath
@@
Appendix I Case Study: Setting alpha=8 instead of alpha=2r
Our main experiments were conducted with $\alpha=2r$. However, Hu et al. (2021) instead set $\alpha=8$ for RoBERTa-base. While not the recommended practice now, we explore what impact this selection has on our findings. We report our key plots in Fig. 14, 15, 16, 17, & 18.
In Fig. 14 & 15 we see that LoRA’d models with high rank have significantly more intruder dimensions in comparison to when $\alpha=2r$. Interestingly, whereas when $\alpha=2r$ LoRA models with ranks like 64 had no or very few intruder dimensions (see Fig. 5), they now have numerous intruder dimensions.
@@ Appendix I Case Study: Setting alpha=8 instead of alpha=2r
These differences are corroborated by Fig. 18, where we see that the learned solutions of LoRA have significantly lower effective rank in comparison to when $\alpha=2r$. For example, we see in Fig.  18 that when LoRA has a rank of 768, the effective rank is never above 100. In contrast, we see in Fig.  13 that with the same rank of 768, LoRA always has an effective rank above 768. This suggests that when $\alpha=8$, LoRA is converging to lower rank solutions than when $\alpha=2r$. This supports the finding that setting $\alpha=2r$ improves LoRA’s performance when a high rank is used (Biderman et al., 2024; Kalajdzievski, 2023).
Behaviorally, we see in Fig. 17 that LoRA models with high rank have much more forgetting on previously learned tasks in comparison to full fine-tuning and LoRA when $\alpha=2r$ is used ($\alpha=2r$ results are in Fig. 8). Likewise, in Fig. 18 we see that when LoRA has high rank, it has much more forgetting on the pre-trained distribution in comparison to LoRA when $\alpha=2r$.

### v2

@@ A.3 Differences in update rule
As described in Appendix E, LoRA and full fine-tuning have characteristically different update rules, even for the same training examples. We highlight that LoRA has gradients projected into a low-rank space [Hao et al., 2024], leading to conditions similar to the toy example in section A.2 above.
A.4 Impact of learning rate on intruder dimensions.
We study the impact learning rate has on intruder dimensions by sweeping a range of learning rates while keeping all other hyperparameters fixed for LoRA $r=8$ and fine-tune on MNLI. Across training epochs, we report the number of intruder dimensions, test accuracy, and pre-training loss (Fig. 7). Across models, we see they have similar test accuracies after 5 epochs but very different numbers of intruder dimensions. We note that small learning rates do not converge as fast as the ones we tested and have difficulty reaching the maximum performance that larger learning rates are able to reach. We see that as we increase learning rate, the number of intruder dimensions also increases (left, Fig. 7). This illustrates a tradeoff in the selection of learning rate of LoRA: picking a larger learning rate may lead to faster convergence and potentially better test accuracy with more intruder dimensions and drift in overall language modeling performance, while smaller learning rates may lead to less drift but potentially lower test accuracy.
Because of this experiment, one may dispute our findings with the claim that they are due to our specific selection of hyperparameters. Therefore, we find it important to note that we adopt hyperparameters from prior literature [Hu et al., 2021], default settings in common machine learning libraries [^2] PEFT, the most popular LoRA library, use learning rates $\geq\mathrm{1e\text{-}3}$ in their tutorials and states “With LoRA-like methods, you can afford to use a higher batch size and learning rate.” [Mangrulkar et al., 2022]., and also study externally trained open-sourced models. This means that our findings are a reflection of common practices and not due to a selection bias by us. Learning rates used by prior work were likely determined based on a variety of factors like speed of convergence and best resulting test accuracy and therefore selected large learning rates that still converge well but coincidentally result in intruder dimensions.
A.5 Matrix product parameterization of LoRA
@@
Figure 10: Impact of only tuning B on the number of intruder dimensions. We randomly initialize A such that it has singular values of 1, freeze it, and only train B. When we do this, we see a sharp reduction in high ranking intruder dimensions in comparison to those in normal LoRA (reported in Fig. 4(d)). Graphs for a specific dataset have the same range as Fig. 4(d) for easy comparison.
@@
Multiplying matrices together amplifies their spectral differences (their singular values) and in most cases leads to a lower effective rank. To test the impact of the product $BA$ on the introduction of intruder dimensions, we randomly initialize $A$ such that all its singular values are 1 and freeze it. We only tune $B$ and keep the rest of our fine-tuning recipe the same. Comparing this with vanilla LoRA is fair because Zhu et al. [2024] found that tuning $B$ is more impactful and important for generalization in comparison to $A$ and Hao et al. [2024] showed that only tuning $B$ effectively approximates LoRA. As we can see in Fig. 10, we see a sharp drop in the number of high ranking intruder dimensions when only tuning $B$ in comparison to the vanilla LoRA case where we train $A$ and $B$ separately, as reported in Fig. 4. This suggests that the matrix product of LoRA is an important component in the introduction of intruder dimensions because of how it amplifies the spectral differences of $B$ and $A$.
Appendix B Implementation Details
B.1 Evaluation details
We follow the precedence of Biderman et al. [2024] when evaluating LLaMA2-7B: When fine-tuned on code, we evaluate on HumanEval [Chen et al., 2021] using the bigcode-eval-harness [Ben Allal et al., 2022]. When fine-tuned on math, we evaluate on GSM8K [Cobbe et al., 2021] using the lm-eval-harness [Gao et al., 2024]. On both, we evaluate task forgetting by evaluating on Hellaswag, WinoGrande, and Arc-Challenge using the lm-eval-harness [Gao et al., 2024].
We measure language modeling loss for all our LLaMA2-7B models on a random sample of its pre-training data distribution, according to Touvron et al. [2023b]. We measure “pseudo-loss” for all our fine-tuned RoBERTa models on a random sample of the four datasets that RoBERTa used for pre-training(OpenWebText [Gokaslan and Cohen, 2019], CCNews [Hamborg et al., 2017], Stories [Trinh and Le, 2019], and bookcorpus [Zhu et al., 2015]) and weigh them proportionally to their contribution as described by Liu et al. [2019].
B.2 Compute Resources
All experiments were run on an internal, shared 8xA100-SXM4-80GB machine. All RoBERTa-base fine-tuning runs required a single A100 GPU. All evaluations and analyses also required a single A100 GPU. Many experiments were run sequentially due to need to share these computing resources. Due to these constraints, instead of fine-tuning our own LLaMA2-7B models, we use publicly released fine-tuned models. For more information on these models, see Section K. Each RoBERTa-base fine-tune run takes at most 6 hours on a single GPU. Evaluating an arbitrary LLaMA2-7B model for both test accuracy and forgetting takes about 45 minutes on a single GPU.
B.3 RoBERTa fine-tuning details
@@
We generally follow the procedure used by Hu et al. [2021]. For all models, we use a linear learning rate schedule with 0.06 linear warmup ratio and train for a maximum of 5 epochs with batch size 16. We use the Adam optimizer [Kingma and Ba, 2017] with no weight decay and a maximum sequence length of 512. We fine-tune all linear layers besides the embedding matrix. For full fine-tuning, we use a learning rate of 1e-5. For LoRA, we set $\alpha=2r$, and train for all ranks in {1, 2, 4, 8, 16, 64}. We hold the “total learning rate of LoRA", which is $\alpha*\eta$, fixed as we sweep rank such that this product always equals 2.4e-3. We fine-tune these models to equivalent accuracy on their downstream task. We fine-tune on six sequence classification tasks: sentiment analysis [Socher et al., 2013], entailment [Williams et al., 2018], duplicate identification [Wang et al., 2019], fact verification [Thorne et al., 2018], and common sense reasoning [Sap et al., 2019, Sakaguchi et al., 2021].
Appendix C Model Accuracies
@@
We report the accuracies that our RoBERTa models achieve in Table 1 and Table 2. Our main results are based on the models in Table 2.
@@
RoBERTa-base Full 0.8745 0.9438 0.9152 0.6582 0.6499 0.6892
r=1 0.8677 0.9415 0.9042 0.6275 0.6418 0.687
r=2 0.869 0.945 0.9054 0.6551 0.6438 0.6822
r=4 0.8698 0.9472 0.9089 0.6361 0.6602 0.6827
r=8 0.8704 0.9472 0.9093 0.6346 0.6607 0.6928
r=16 0.8739 0.9461 0.9093 0.6417 0.6571 0.6924
r=64 0.8719 0.9472 0.9061 0.6212 0.6167 0.6864
Table 2: Model accuracies on their given downstream task after fine-tuning for $\alpha=2r$. Our main results are based on these models.
Appendix D Cosine Similarity with Orthogonal Vectors that Span a Space
@@
Appendix E Derivation of LoRA Adapter’s Gradients
Our calculations were derived independently but follow a similar line to that of Hao et al. [2024].
@@ Appendix E Derivation of LoRA Adapter’s Gradients
Appendix F Intruder Dimensions Correlate with Forgetting
F.1 For RoBERTa
As mentioned in the main text, when measuring the Spearman correlation between the number of intruder dimensions and forgetting in Fig. 7 we find an extremely strong fit, with $\rho=0.971$ and p-value $\ll 0.001$. This shows us that intruder dimensions strongly correlate with forgetting. In contrast, when we correlate intruder dimensions with performance, we find no such correlation: for RoBERTa, we measure $\rho=-0.3381$ with p-value $=0.218$.
F.2 LLaMA2-7B
When measuring the Spearman correlation between number of intruder dimensions and forgetting for our LLaMA2-7B models across training epochs (Fig. 11 (middle)), we find a statistically significant relationship with $\rho=0.59$ and p-value $=0.0006$. When correlating intruder dimensions and test accuracy (Fig. 11 (left)), we instead measure $\rho=-0.3178$ with p-value $=0.0869$. Again, we see that intruder dimensions correlates with forgetting.
Figure 11: For LLaMA2-7B, intruder dimensions correlate with forgetting. Top row: MetaMath. Bottom row: Magicoder. We display intruder dimensions vs test accuracy and intruder dimensions vs forgetting.
Appendix G Intruder Dimensions Cause Forgetting (Scaling Experiments)
G.1 Performance Differences When Scaling Down Intruder Dimensions
G.1.1 RoBERTa
We report our findings for RoBERTa models fine-tuned on MNLI, QQP, and FEVER in Table 3. Remember that we scale down using the equation $W=W_{0}+\Delta W+(\lambda-1)u_{i}\sigma_{i}v^{T}_{i}$. Here, we see that scaling down intruder dimensions leads to a sharp drop in forgetting (pre-training loss) and a much smaller drop in test accuracy. Scaling down an intruder dimension by two ($\lambda=0.5$) results always leads to less than a two percent drop in test accuracy but double digit percentage drops in forgetting. One particularly compelling example, as shown in Fig. 8(d), shows how scaling down the top intruder dimensions with $\lambda=0.7$ when fine-tuning on QQP and using LoRA r=8 result in essentially no (0.0%) drop in adaptation performance but a large (-33.2%) drop in forgetting. Our findings of the impact of scaling down intruder dimensions hold across three datasets and both LoRA r=1 and r=8, which were the two ranks that we found to have many intruder dimensions. Note that this experiment is meaningless if a model has no intruder dimensions, since no singular vectors will be removed.
Task LoRA $\lambda=0.1$ $\lambda=0.3$ $\lambda=0.5$ $\lambda=0.7$ $\lambda=0.9$
Rank TA PTL TA PTL TA PTL TA PTL TA PTL
MNLI r=1 -18.7 -13.3 -8.3 -14.5 -2.7 -13.7 -0.6 -11.0 0.0 -5.1
r=8 -5.1 -24.7 -1.9 -23.1 -0.6 -19.8 -0.2 -14.5 0.0 -5.9
QQP r=1 -8.6 -35.5 -4.4 -35.8 -1.6 -34.1 -0.4 -28.9 0.1 -14.6
r=8 -3.3 -52.0 -1.6 -50.6 -0.6 -45.0 -0.0 -33.2 0.1 -13.0
FEVER r=1 -11.1 -10.3 -4.2 -11.8 -0.4 -11.1 0.6 -8.7 0.6 -3.8
r=8 -5.6 -14.6 -1.0 -15.3 0.7 -13.4 1.3 -9.7 0.6 -4.0
Table 3: Impact of scaling RoBERTa-base’s intruder dimensions on test accuracy (TA) and pre training loss (PTL). Numbers reported are the percent change in test accuracy and percent reduction in forgetting induced by fine-tuning. Scaling down intruder dimensions leads to less forgetting. On RoBERTa. PTPL is Pre-training loss and TA is test accuracy. Both are reported as percent change with respect to the unchanged fine-tuned model. Scaling down intruder dimensions has large impact on forgetting but little impact on test accuracy.
G.1.2 LLaMA2-7B
Our finding that scaling down intruder dimensions leads to less forgetting but similar test accuracy holds to LLaMA2-7B. One particularly interesting example is for LLaMA2-7B fine-tuned on MetaMath with $r=256$: when scaling the top intruder dimensions down with $\lambda=0.5$, we see a large drop (-25.2%) in forgetting and an increase (+1.8%) in test accuracy.
Dataset LoRA $\lambda=0.3$ $\lambda=0.5$ $\lambda=0.7$ $\lambda=0.9$
Rank TA PTL TA PTL TA PTL TA PTL
MetaMath r=16 -15.0 -46.5 -6.6 -40.4 -2.7 -28.2 0.0 -10.5
r=64 -5.8 -29.1 -4.4 -22.3 0.3 -14.2 0.3 -4.9
r=256 -0.1 -33.3 1.8 -25.2 0.5 -15.5 -0.8 -5.2
Magicoder r=16 -1.6 -37.5 -0.6 -24.4 -0.1 -13.2 0.1 -4.1
r=64 -5.0 -12.7 -3.3 -8.3 0.3 -4.3 2.2 -1.4
r=256 -4.3 -12.8 -1.3 -10.2 -0.9 -7.0 -1.6 -2.8
Table 4: Impact of scaling LLaMA2-7B’s intruder dimensions on test accuracy (TA) and pre training loss (PTL). Numbers reported are the percent change in test accuracy and percent reduction in forgetting induced by fine-tuning.
G.2 Intruder Dimensions Cause Worse OOD Performance
As we discussed in section 5 of the main text, for LoRA models with main intruder dimensions ($r=1$ and $r=8$ in our experiments) we measure the impact of intruder dimensions by identifying and scaling the top intruder dimension in every weight matrix such that $W=W_{0}+\Delta W+(\lambda-1)u_{i}\sigma_{i}v^{T}_{i}$, where $i$ is the index of the top intruder dimension (note that $\lambda=0$ is removal, $\lambda=1$ is no change, and $\lambda=2$ doubles the intruder dimension). For our RoBERTa models fine-tuned on MNLI, QQP, and FEVER, we use $\lambda\in\{0,0.25,0.5,0.75,1.0,1.5,2.0\}$ and for each measure the test accuracy and pre-training loss. For a comparison baseline, we select the neighbor of the intruder dimension to separately scale.
We report these results in Fig. 12(a) (MNLI), Fig. 12(b) (QQP), and Fig. 12(c) (FEVER). We find that when scaling down intruder dimensions ($\lambda<1$), we observe a clear and significant drop in forgetting (pre-training loss) but a negligible drop in test accuracy (adaptation). When we instead scale up intruder dimensions ($\lambda>1$), we observe that forgetting increases significantly. We observe that when scaling up top intruder dimensions by 50% ($\lambda=1.5$), adaptation performance remains relatively flat with a large increase in forgetting, providing further evidence that intruder dimensions are task specific. These trends hold across all 6 models we study (3 datasets with two different LoRA ranks).
In contrast, when we scale a neighboring (pre-trained) singular vector of intruder dimensions instead, we observe starkly different behaviors. When scaling down ($\lambda<1$) these pre-trained singular vectors, we observe we see that forgetting sharply increases and adaptation performance drops more sharply than when scaling intruders instead. When scaling up ($\lambda>1$) these pre-trained singular vectors, we observe similar drops in forgetting and adaptation performance. This is likely because pre-trained singular vectors are well tuned for language modeling, and therefore any change to them will have negative downstream impacts on performance. This shows that are observations are not due to the robustness of the model to scaling down specific singular vectors, but rather the difference in contribution to model performance of intruder dimensions vs. pre-trained singular vectors.
Here, we clarify some possible points of confusion. We only scale the top intruder dimension in each weight matrix, if and only if an intruder exists in that matrix, so even when $\lambda=0$ we do not recover the pre-trained model because not all weight matrices have intruder dimensions and it is possible for multiple intruder dimensions to exist in a weight matrix if LoRA $r>1$. Furthermore, intruder dimensions are not perfectly orthogonal to the pre-trained singular vectors. Due to the orthogonality constraint imposed by the SVD, all singular vectors will be changed slightly in the matrix, so that even when we remove the intruder dimension, the resulting matrix will be slightly different. These reasons are why when $\lambda=0$ our LoRA $r=1$ models do not return to baseline performance.
These findings hold to LLaMA2-7B: in Fig 13, we see that scaling down intruder dimensions leads to much less forgetting but similar test accuracy.
(a) RoBERTa-base fine-tuned on MNLI.
(b) RoBERTa-base fine-tuned on QQP.
(c) RoBERTa-base fine-tuned on FEVER.
Figure 12: Scaling RoBERTa-base’s intruder dimensions. We scale the top intruder dimension in each matrix by $\lambda$, a multiplicative constant, such that $W=W_{0}+\Delta W+(\lambda-1)u_{i}\sigma_{i}v^{T}_{i}$. Using $\lambda<1$ leads to a large drop in pre-training loss while only slightly impacting the test accuracy. For Figs. 12(a), 12(b), and 12(c), we also scale the intruder dimension’s neighbor, which is a pre-trained singular vector. Changing these vectors negatively impacts both pre-training loss and test accuracy.
Figure 13: Scaling down LLaMA2-7B’s intruder dimensions leads to less forgetting and nearly equivalent test accuracy.
Appendix H The Effective Rank of the Update Matrix Depends on Alpha
Kalajdzievski [2023] found that LoRA can have gradient collapse when rank is high if alpha is not set properly. Biderman et al. [2024] found that setting $\alpha=2r$ is very important for the performance of high rank LoRA. In this section, we provide additional evidence of the importance of setting $\alpha=2r$ and show that if $\alpha$ is held fixed, high rank LoRA converges to low rank solutions. To do this, for both $\alpha=2r$ and $\alpha=8$, we measure the effective rank [Roy and Vetterli, 2007] of the weight matrix updates for different LoRA ranks. Effective rank is a measure of the information density of a matrix and can be thought of as an estimation of the rank needed to capture the information held in the weight matrix. It is computed using the singular values of a matrix and we expect the LoRA rank to be the upper bound on what the effective rank can be: LoRA $r=8$ should have an effective rank update of at most 8. We present the effective rank measurements in Fig. 14. In this plot, we find that when $\alpha=2r$ (Fig. 14(a)), LoRA $r=64$ and $r=768$ have much higher effective rank than $r=8$ and $r16$, with $r=768$ appearing to always have effective rank above 100. In stark contrast, we see that when $\alpha=8$ (Fig. 14(b)), high ranks of LoRA have much lower effective ranks, frequently even converging to the effective rank of much lower rank updates (like $r=8$ and $r=16$). For example, LoRA $r=768$ has an effective rank that is consistently below 50 when $\alpha=8$. We note that full fine-tuning has an effective rank update of above 400 consistently. These plots suggest that when $\alpha$ is kept fixed when scaling LoRA rank, the solutions are uable to take advantage of their higher expressability and instead converge to low rank solutions.
(a) Effective Rank of the LoRA update when $\alpha=2r$.
(b) Effective Rank of the LoRA update when $\alpha=8$.
Figure 14: Effective rank of LoRA update matrices ($\Delta W$) for RoBERTa fine-tuned on MNLI. We observe that when $\alpha=2r$, higher ranks of LoRA ($r=64,768$) have much higher effective rank than the same ranks of LoRA but instead with $\alpha=8$. Building on Kalajdzievski [2023], Biderman et al. [2024], this suggests that $\alpha=2r$ is necessary for high ranks of LoRA to utilize their expressive capacity. Note: full fine-tuning consistently has updates with effective rank above 400.
Appendix I Impact of Matrix Percentage on Number of Intruder Dimensions
In this section, we examine the extent to which intruder dimensions exist throughout the entire weight matrix and how they are distributed. As described in the main text, we hold $\epsilon$ fixed as $\epsilon=0.5$ and measure the number of intruder dimensions while varying the proportion of the fine-tuned singular vectors that we examine (this means varying our $k$ parameter in Algorithm 1). Here, we can see that LoRA consistently has more intruder dimensions than full fine-tuning, regardless of what fraction of the singular values we examine. The only caveat to this is that, for some datasets, full fine-tuning passes LoRA $r=1$ when examining the last 20% of the fine-tuned singular vectors. This is likely due to the limited expressivity of rank 1 updates and is interesting because it suggests that in this case, full fine-tuning may be changing lower-ranking singular vectors more than LoRA. One interesting contradiction to our findings is in Fig. 15(d), which shows that full fine-tuning and LoRA appear to have very similar distributions of intruder dimensions within their matrix when fine-tuned on code. This is likely due to the large domain shift from natural language to coding tasks (Biderman et al. [2024] also make this observation of a large domain shift required for models fine-tuned on Magicoder [Wei et al., 2024]).
@@ Appendix I Impact of Matrix Percentage on Number of Intruder Dimensions
Appendix J Impact of Dataset Size On Intruder Dimensions
The total number of intruder dimensions increases proportionally to the size of the fine-tuning dataset. Using our training recipe (Appendix B.3), we fine-tuned models on data subsets of varying sizes. We trained RoBERTa-base on MNLI using LoRA with rank 1 and 8 (cases where we originally saw intruder dimensions) and measure the number of intruder dimensions along with the impact of $\epsilon$ and $k$ (Fig. 16). For $r=8$, as we train on more data, more intruder dimensions are introduced. Interestingly, however, LoRA with rank 1 appears to converge to similar amounts of intruder dimensions, regardless of the dataset size. This may be because of the limited expressivity of models with $r=1$. This experiments suggest that with smaller datasets, fewer intruder dimensions may be introduced by LoRA.
Figure 16: (Left) Impact of cosine similarity threshold, $\epsilon$, on the number of intruder dimensions for LoRA models trained on different proportions of the MNLI dataset. (Right) Impact of the number of fine-tuned singular vectors we examine, $k$, on the number of intruder dimensions for LoRA models trained on different proportions of the MNLI dataset. We see that training on a larger proportion of the dataset increases the number of intruder dimensions in the model.
Appendix K LLaMA/LLaMA-2 Instruction Tuned Models
Our LLaMA-7B checkpoints were fine-tuned on the Alpaca [Taori et al., 2023] and consist of two fully fine-tuned models, one LoRA model with rank 16, and one QLoRA [Dettmers et al., 2023] model with rank 64. Our LLaMA2-7B checkpoints were fine-tuned on either code (IFT with Magicoder-Evol-Instruct-110K [Wei et al., 2024] or CPT with StarCoder [Li et al., 2023]) or math (IFT with MetaMathQA [Yu et al., 2024] or CPT with OpenWebMath [Paster et al., 2023]) and consist of one fully fine-tuned model and 3-4 LoRA’ed models of different ranks for each dataset and generously provided by Biderman et al. [2024]. In Fig. 4(a), Full #1 refers to “PKU-Alignment/alpaca-7b-reproduced" and Full #2 refers to “chavinlo/alpaca-native".
@@
LoRA-TMLR-2024/metamath-lora-rank-16-alpha-32 LLaMA2-7b MetaMath
LoRA-TMLR-2024/metamath-lora-rank-64-alpha-128 LLaMA2-7b MetaMath
LoRA-TMLR-2024/metamath-lora-rank-256-alpha-512 LLaMA2-7b MetaMath
LoRA-TMLR-2024/metamath-full-finetuning-lr-1e-05 LLaMA2-7b MetaMath
LoRA-TMLR-2024/starcoder-lora-rank-16-20B-tokens LLaMA2-7b StarCoder
LoRA-TMLR-2024/starcoder-lora-rank-64-20B-tokens LLaMA2-7b StarCoder
LoRA-TMLR-2024/starcoder-lora-rank-256-20B-tokens LLaMA2-7b StarCoder
LoRA-TMLR-2024/starcoder-full-finetuning-lr-1e-05-20B-token LLaMA2-7b StarCoder
LoRA-TMLR-2024/openwebmath-lora-rank-16-20B-tokens LLaMA2-7b OpenWebMath
LoRA-TMLR-2024/openwebmath-lora-rank-64-20B-tokens LLaMA2-7b OpenWebMath
LoRA-TMLR-2024/openwebmath-lora-rank-256-20B-tokens LLaMA2-7b OpenWebMath
LoRA-TMLR-2024/openwebmath-full-finetuning-lr-1e-05-20B-tokens LLaMA2-7b OpenWebMath
@@
Appendix L Continual Learning
L.1 Performance during continual learning
As described in Section 4 in the main text, we train sequentially on 6 tasks and measure task performance on all of these tasks across tasks trained on (continual learning). We report the full graph of our findings in Fig. 17. In it, we find that when all our models are trained to similar accuracy, lower ranks of LoRA, which coincide with more intruder dimensions, forget more of their previously learned tasks than higher ranks of LoRA and full fine-tuning.
Figure 17: Full plot of Fig. 9(a). Continual Learning performance of RoBERTa for full fine-tuning and LoRA. We sequentially train on six tasks, in order from left to right. Horizontal dotted line indicates baseline pre-trained performance. Vertical solid line indicates when a specific dataset is fine-tuned on. Gray region represents performance before the model has been trained on that task. We are interested in the differences in accuracies of these methods both right after training (at the vertical black line) and later (in the white region). We see that low ranks of LoRA forget previously learned tasks more.
L.2 Similarity matrices during continual learning
After each continual learning dataset we fine-tune on, we measure the similarity matrix between the current model and the pre-trained model. In Fig. 18(b), we observe that LoRA accumulates intruder dimensions across fine-tuning datasets. In contrast, in Fig. 18(a) we observe that the pre-trained structure of the model is retained well across fine-tuning datasets. These experiments suggest why LoRA appears to degrade faster during continual learning.
(a) Continual learning similarity matrices for full fine-tuning.
(b) Continual learning similarity matrices for LoRA.
Figure 18: LoRA accumulates intruder dimensions, while full fine-tuning does not. The pre-trained structure of the model degrades across tasks trained on.
Appendix M Case Study: Setting Alpha=8 instead of Alpha=2r
@@
Our main experiments were conducted with $\alpha=2r$. However, Hu et al. [2021] instead set $\alpha=8$ for RoBERTa-base. While not the recommended practice now, we explore what impact this selection has on our findings. We report our key plots in Fig. 19(a), 19(b), 20, 21, & 14(b). In Fig. 19(a) & 19(b) we see that LoRA’d models with high rank have significantly more intruder dimensions in comparison to when $\alpha=2r$. Interestingly, whereas when $\alpha=2r$ LoRA models with ranks like 64 had no or very few intruder dimensions (see Fig. 4), they now have numerous intruder dimensions. These differences are corroborated by Fig. 14(b), where we see that the learned solutions of LoRA have significantly lower effective rank in comparison to when $\alpha=2r$. For example, we see in Fig.  14(b) that when LoRA has a rank of 768, the effective rank is never above 100. In contrast, we see in Fig.  14(a) that with the same rank of 768, LoRA always has an effective rank above 768. This suggests that when $\alpha=8$, LoRA is converging to lower rank solutions than when $\alpha=2r$. This supports the finding that setting $\alpha=2r$ improves LoRA’s performance when a high rank is used [Biderman et al., 2024, Kalajdzievski, 2023]. Behaviorally, we see in Fig. 21 that LoRA models with high rank have much more forgetting on previously learned tasks in comparison to full fine-tuning and LoRA when $\alpha=2r$ is used ($\alpha=2r$ results are in Fig. 17). Likewise, in Fig. 14(b) we see that when LoRA has high rank, it has much more forgetting on the pre-trained distribution in comparison to LoRA when $\alpha=2r$.